# JoPA: Explaining Large Language Model’s Generation via Joint Prompt Attribution

### v2

@@ 1 Introduction
Large Language Models (LLMs), such as GPT4 (achiam2023gpt, ), LLaMA (touvron2023llama, ) and Claude (anthropic_2024_claude, ), have shown excellent performance in various natural language generation tasks including question answering, document summarization, and many more. Despite the great success of LLMs, we still have very limited understanding of the LLM generation behavior – which parts in the input cause the model to generate a certain sequence. Unable to explain the causality between the input prompt and the output generation could cause failure in recognizing potential unintended consequences, such as harmful response (dan, ; autodan-liu, ; gcg, ; autodan-zhu, ) and biased generation (wang2023decodingtrust, ) attributed to a specific malicious description in the input. These issues undermine human trust in model usage, thus highlighting a pressing need for developing an interpretation tool that attributes how an input prompt leads to the generated content.
@@
Explaining LLM generation through prompt attribution involves extracting the most influential prompt texts on the model’s entire generation procedure, a realm that remains relatively under-explored in current research. While extensive works on input attribution are proposed for text classification interpretation (chen2020learning, ; modarressi2023decompx, ) and next-word generation rationale (zhao2024reagent, ; vafa2021rationales, ), they can not be directly applied to explain the full generation sequence due to its complicated joint probability landscape and the autoregressive generation procedure. The complexity for interpreting LLM generation compounds as the model size increases. Another line of works involves prompting LLMs to self-explain their behaviors (wei2022chain, ). This method relies on the model’s innate reasoning capabilities, although current findings suggest that these capabilities may not always be faithful (Turpin2023LanguageMD, ; xu2024hallucination, ).
@@
There are limited existing attempts focusing on explaining the relationship between the input prompts and the complete generated sequence. The most relevant approach, Captum (captum, ), sequentially determines the importance score for each token by calculating the variations in the joint probability of generating the targeted output sequence when the token is dropped from the model input. While being straightforward, this approach treats tokens as independent features, ignoring their joint semantic influence on the generated output. In fact, tokens may contain overlapping or complementary information. For example, given the input prompt: “Write a story about the doctor and his patient”, the most influential components are “doctor” and “patient”. Individually removing either of these words would not significantly alter the generated output, resulting in inaccurately low importance score for each of them. This is caused by the semantic interaction among these components, allowing the model to infer the meaning of the omitted one. Existing attribution methods that ablate each token individually fail to capture such combinatorial effect. A straightforward remedy might involve exhaustively assessing all possible combinations to observe the variations in the model generations, which however is impractical due to the vast search space with long-context input prompts.
@@ 2 Related Work
While there are extensive works devoted to explaining language models in the context of text classification tasks (shi2022nearest, ; han2021ptr, ; shi2023chatgraph, ), relatively few attempts (zhao2024reagent, ; vafa2021rationales, ) are proposed to investigate the importance of prompt texts on the entire generation procedure especially for LLMs. This demonstrates a research gap that this work aims to fill in.
@@ 2.1 Explaining Language Generation
There are many work explaining the predictions generated by LLMs by measuring the importance of the input features to the model’s prediction on the classification tasks. One group of studies perturb the specific input by removing, masking, or altering the input features, and evaluate the model prediction changes (10.1145/3461702.3462597, ; wu-etal-2020-perturbed, ). The other group of works, such as integrated gradients (IG) (sundararajan2017axiomatic, ), first-derivative saliency (li-etal-2016-visualizing, ), and mixed partial derivatives (tsang2020does, ) leverage the gradients of the output with respect to the input to determine the input feature importance. Although Archipelago (tsang2020does, ) explains the feature attributions by considering the combined effects of the input attributions, it targets the multilabel classification task and relies on the neutral baseline. As for the generation tasks, Diffmask learns the differentiable mask for each layer of the BERT model, but it aims to analyze how decisions are formed across network hidden layers by a simple probing classifier. In contrast, our framework targets on offering insights into the relationship between the input prompt and model generations by employing the mask to highlight the essential prompt attributions.
Moreover, a few studies utilize the surrogate model to explain the individual predictions of the black-box models, and the representative method is called LIME (ribeiro2016should, ). As for explaining the model’s generation behavior, Captum (captum, ) calculates the token’s importance score by sequentially measuring the contribution of input token to the output, which lacks accounting for the semantic relationships between tokens. Another work, ReAGent (zhao2024reagent, ), focuses on the next-word generation task, computing the importance distribution for the next token position. This method ignores the contextual dependencies in the generated output and could not adequately account for dynamically changing generations. Our framework aims to interpret the joint effects of the input prompts on the entire output contexts with considering of the textual information covered the input prompt.
@@ 2.2 Self-explaining by Prompting
As language models increase in scale, prompting-based models demonstrate remarkable abilities in reasoning (brown2020language, ), creativity (oppenlaender2022creativity, ), and adaptability across a range of tasks (khattak2023maple, ). However, the complex reasoning processes of these models remain elusive and require tailored paradigm to better understand the prompting mechanism. For instance, the chain-of-thought (CoT) paradigm could explain the LLM behaviors by prompting the model to generate the reasoning chain along with the answers (wei2022chain, ), as pre-trained LLMs have demonstrated a certain ability to self-explain their behaviors. However, recent studies have also suggested that the reasoning chain does not guarantee faithful explanations of the model’s behavior (jacovi-goldberg-2020-towards, ) and the final answer might not always follow the generated reasoning chain (Turpin2023LanguageMD, ). xLLM (chuang2024large, ) improves explanation fidelity from LLMs using a fidelity evaluator, tailored for classification tasks. We focus on explaining LLM generation by analyzing input prompt attributions to output content, independent of the model’s suboptimal reasoning abilities.
@@ 3 Preliminaries
There are limited prior works about explaining the relationship between the individual input prompts and the entire generated sequence, and their faithfulness is limited by treating input tokens independently. Specifically, Captum (captum, ), calculates the importance score for each token by slightly perturbing the input $\bm{x}$ at the $i$-th token:
@@ 4 Proposed Method
In order to explain language generation via prompt attribution, instead of treating the prompt tokens independently as in previous works (captum, ), we propose to evaluate the joint effect of $k$ prompt tokens on the generated sequence.
@@
Gradient-Guided Masking: Trivial solutions that set the mask $\bm{m}$ by uniformly random could cost massive sampling to hit the right optimization direction. Gradient is a common indicator of feature importance, as evidenced in existing practices (ebrahimi2018hotflip, ; gcg, ; shin-etal-2020-autoprompt, ). Therefore, we propose to use gradient as a guidance to set and sample the mask for more efficient optimization. Specifically, we begin with the binary mask $\bm{m}^{(0)}=\bm{1}$ which indicates that all tokens in the input $\bm{x}$ are marked as non-explanatory ones. Then we compute the gradient $\nabla_{\bm{m}^{(0)}}\mathcal{L}(\bm{m}^{(0)},\bm{x},\bm{y};\bm{\theta})$ of the loss function in Eq. (3), and denote the magnitude of gradients as $\bm{g}=|\nabla_{\bm{m}^{(0)}}\mathcal{L}(\bm{m}^{(0)},\bm{x},\bm{y};\bm{\theta})|$. Note that the components with larger gradient magnitudes in $\bm{g}$ imply that altering the corresponding tokens could result in a sharper change to the generated outputs. In order to initialize the explanatory $k$ tokens guided by the gradient magnitude, we set the binary mask $\bm{m}^{(1)}$ where $m^{(1)}_{i}=0$ indicates $g_{i}$ is the top-$k$ value in $\bm{g}$. The gradient guidance and mask initialization are obtained following Line 1-2 in the algorithm.
@@ 5.1 Experiment Settings
In the experiment, we employ two LLMs as our targeted $f_{\bm{\theta}}(\cdot)$: LlaMA-2 (7B-Chat) (touvron2023llama, ) and Vicuna (7B) (Zheng2023JudgingLW, ). There are relatively few methods attributing prompts on the entire language generation, and we choose random removal (Random), Integrated-Gradient (sundararajan2017axiomatic, ), averaged attentions across all layers (pruthi2019learning, )(Attention), last-layer attention (zhao2024reagent, )(Last-Attention) and Captum (captum, ) as our baseline for comparison. We also compare JoPA with ReAGent (zhao2024reagent, ), with the results presented in Appendix A.13. We implement these models using the PyTorch framework and pretrained weights from the transformers Python library (wolf-etal-2020-transformers, ), and conduct our experiments on an Nvidia RTX A6000-48GB platform with CUDA version 12.0.
@@
We employ three distinct text generation datasets: Alpaca (alpaca, ), tldr_news (tldr_news, ), and MHC (amod_2024, ), to evaluate the effectiveness of our method across various generation tasks. To capture the joint influence of prompts on model outputs, we use long-context prompts (15+ words) rather than simple sentences, as they convey more information and exhibit higher textual correlation. For evaluation, 110 samples are randomly selected from each publicly available dataset, detailed in AppendixA.2
@@ 5.2 Evaluation Metrics
Faithfulness scores are a key metric for assessing the quality of explanations, with faithful explanations accurately reflecting the model’s decision-making process (jacovi-goldberg-2020-towards, ). Studies (7552539, ; hooker2019benchmark, ) suggest that if a certain input tokens are truly important, their removal should lead to a more significant change in model output than the removal of random tokens (madsen-etal-2022-evaluating, ). Therefore, after removing explanatory tokens identified by a faithful method, the model’s new generation would significantly differ from using the original output. Moreover, the input prompt with masking these explanatory tokens are less likely to reproduce the original model response $\bm{y}$, indicating a lower value of $p_{\bm{\theta}}(\bm{y}|\bm{m}\odot\bm{x})$.
@@
Word Frequency: BLEU (bleu, ) evaluates how closely the candidate text $\bm{y}^{\prime}$ matches the reference $\bm{y}$ by measuring n-gram precision with clipping to avoid overcounting and brevity adjustments for shorter candidates. ROUGE-L (rouge, ) measures sequence overlap between $\bm{y}^{\prime}$ and $\bm{y}$, assessing precision, recall, and their F1 score.
Semantic Similarity: SentenceBert (reimers2019sentencebert, ) is a variation of the BERT model which is designed to generate high-quality sentence embeddings for the pairs of sentences. We leverage SentenceBert to transform the text $\bm{y}$ and $\bm{y}^{\prime}$ into fixed-length embedding vectors, and calculate the cosine similarity between their embeddings to quantify their semantic similarity.
Probability Measurement: We define two measurements to reflect how the likelihood of generating the original text $\bm{y}$ changes before and after applying the explanation mask. We first define the Probability Ratio (PR) to indicate how less likely to generate the original output $\bm{y}$ when explanatory tokens are masked: $\text{PR}(\bm{x},\bm{y},\bm{m})=\frac{\tilde{p}_{\bm{\theta}}(\bm{y}|\bm{m}\odot\bm{x})}{\tilde{p}_{\bm{\theta}}(\bm{y}|\bm{x})}$, where $\tilde{p}(\bm{y}|\bm{x})=p(\bm{y}|\bm{x})^{1/s}$ is the length-normalized generation probability. If the PR score is far below the random baseline, we can conclude that the masked tokens are indeed important to cause the model generating $\bm{y}$. In addition, we also calculate the KL-divergence between these two distributions to measure their difference: $D_{\text{KL}}\large(p_{\bm{\theta}}(\bm{y}|\bm{m}\odot\bm{x})||p_{\bm{\theta}}(\bm{y}|\bm{x})\large)$, and a larger score indicates a larger change in generation likelihood and a more accurate explanation.
@@ A.2 Dataset Details
  • Alpaca alpaca is a dataset with 52000 unique examples consisting of instructions and demonstrations generated by OpenAI’s text-davinci-003 engine. Each example in the dataset includes an instruction that describes the task the model should perform, accompanied by optional input context for that task. In our experiment, we randomly select a subset of these examples for verification purposes.
  • tldr_news tldr_news dataset is constructed by collecting a daily tech newsletter. For every piece of data, there is a headline and corresponding content extracted. The task is to ask the model to simplify the extracted content and then generate a headline from the input.
  • mental_health_counseling (MHC) amod_2024 includes broad pairs of questions and answers derived from online counseling and therapy platforms. It covers a wide range of mental-health related questions and concerns, as well as the advice provided by the psychologists. It is utilized for used to fine-tuning the model to generate the metal health advice. Here, we prompt the model to generate an advice based on the provided question.
@@ A.3 Case Study
In this section, we visualize three examples which are selected from the dataset Alpaca, tldr_news, and MHC respectively, shown in Figure. 3. The tokens in the red frame are the $k$ explanatory tokens extracted by the model LlaMA-2 (7B-Chat), and the output responses framed in blue are the most informative results selected by human. In the first data sample, the recognized token: "brains" and "cognitive", has some semantic relationship intuitively. It is a common sense the cognitive ability is underpinned by the brain’s function and structure. There are numerous reports and studies focused on the functionally relationship between the brain and cognitive ability (zhang2019cognitive, ). Due to the extensive amount of data on which the LLMs are trained, the model can integrate their relationship which is reflected in its response.

### v3

@@ Abstract
JoPA: Explaining Large Language Model’s Generation via Joint Prompt Attribution
Yurui Changthanks: Equal Contribution, Bochuan Cao11footnotemark: 1, Yujia Wang, Jinghui Chen, Lu Lin Pennsylvania State University {yuruic, bccao, yjw5427, jzc5917, lulin}@psu.edu
@@ 1 Introduction
Large Language Models (LLMs), such as GPT4 (Achiam et al., 2023), LLaMA (Touvron et al., 2023) and Claude (Anthropic, 2024), have shown excellent performance in various natural language generation tasks including question answering, document summarization, and many more. Despite the great success of LLMs, we still have very limited understanding of the LLM generation behavior – which parts in the input cause the model to generate a certain sequence. Unable to explain the causality between the input prompt and the output generation could cause failure in recognizing potential unintended consequences, such as harmful response (DAN, 2023; Liu et al., 2023; Zou et al., 2023; Zhu et al., 2023) and biased generation (Wang et al., 2023) attributed to a specific malicious description in the input. These issues undermine human trust in model usage, thus highlighting a pressing need for developing an interpretation tool that attributes how an input prompt leads to the generated content.
@@
Explaining LLM generation through prompt attribution involves extracting the most influential prompt texts on the model’s entire generation procedure, a realm that remains relatively under-explored in current research. While extensive works on input attribution are proposed for text classification interpretation (Chen and Ji, 2020; Modarressi et al., 2023) and next-word generation rationale (Zhao and Shan, 2024; Vafa et al., 2021), they can not be directly applied to explain the full generation sequence due to its complicated joint probability landscape and the autoregressive generation procedure. The complexity for interpreting LLM generation compounds as the model size increases. Another line of works involves prompting LLMs to self-explain their behaviors (Wei et al., 2022). This method relies on the model’s innate reasoning capabilities, although current findings suggest that these capabilities may not always be faithful (Turpin et al., 2023; Xu et al., 2024).
@@
There are limited existing attempts focusing on explaining the relationship between the input prompts and the complete generated sequence. The most relevant approach, Captum (Miglani et al., 2023), sequentially determines the importance score for each token by calculating the variations in the joint probability of generating the targeted output sequence when the token is dropped from the model input. While being straightforward, this approach treats tokens as independent features, ignoring their joint semantic influence on the generated output. In fact, tokens may contain overlapping or complementary information. For example, given the input prompt: “Write a story about the doctor and his patient”, the most influential components are “doctor” and “patient”. Individually removing either of these words would not significantly alter the generated output, resulting in inaccurately low importance score for each of them. This is caused by the semantic interaction among these components, allowing the model to infer the meaning of the omitted one. Existing attribution methods that ablate each token individually fail to capture such combinatorial effect. A straightforward remedy might involve exhaustively assessing all possible combinations to observe the variations in the model generations, which however is impractical due to the vast search space with long-context input prompts.
@@ 2 Related Work
While there are extensive works devoted to explaining language models in the context of text classification tasks (Shi et al., 2022; Han et al., 2021; Shi et al., 2023), relatively few attempts (Zhao and Shan, 2024; Vafa et al., 2021) are proposed to investigate the importance of prompt texts on the entire generation procedure especially for LLMs. This demonstrates a research gap that this work aims to fill in.
@@ 2.1 Explaining Language Generation
There are many work explaining the predictions generated by LLMs by measuring the importance of the input features to the model’s prediction on the classification tasks. One group of studies perturb the specific input by removing, masking, or altering the input features, and evaluate the model prediction changes (Kommiya Mothilal et al., 2021; Wu et al., 2020). The other group of works, such as integrated gradients (IG) (Sundararajan et al., 2017), first-derivative saliency (Li et al., 2016), and mixed partial derivatives (Tsang et al., 2020) leverage the gradients of the output with respect to the input to determine the input feature importance. Although Archipelago (Tsang et al., 2020) explains the feature attributions by considering the combined effects of the input attributions, it targets the multilabel classification task and relies on the neutral baseline. As for the generation tasks, Diffmask learns the differentiable mask for each layer of the BERT model, but it aims to analyze how decisions are formed across network hidden layers by a simple probing classifier. In contrast, our framework targets on offering insights into the relationship between the input prompt and model generations by employing the mask to highlight the essential prompt attributions.
Moreover, a few studies utilize the surrogate model to explain the individual predictions of the black-box models, and the representative method is called LIME (Ribeiro et al., 2016). As for explaining the model’s generation behavior, Captum (Miglani et al., 2023) calculates the token’s importance score by sequentially measuring the contribution of input token to the output, which lacks accounting for the semantic relationships between tokens. Another work, ReAGent (Zhao and Shan, 2024), focuses on the next-word generation task, computing the importance distribution for the next token position. This method ignores the contextual dependencies in the generated output and could not adequately account for dynamically changing generations. Our framework aims to interpret the joint effects of the input prompts on the entire output contexts with considering of the textual information covered the input prompt.
@@ 2.2 Self-explaining by Prompting
As language models increase in scale, prompting-based models demonstrate remarkable abilities in reasoning (Brown et al., 2020), creativity (Oppenlaender, 2022), and adaptability across a range of tasks (Khattak et al., 2023). However, the complex reasoning processes of these models remain elusive and require tailored paradigm to better understand the prompting mechanism. For instance, the chain-of-thought (CoT) paradigm could explain the LLM behaviors by prompting the model to generate the reasoning chain along with the answers (Wei et al., 2022), as pre-trained LLMs have demonstrated a certain ability to self-explain their behaviors. However, recent studies have also suggested that the reasoning chain does not guarantee faithful explanations of the model’s behavior (Jacovi and Goldberg, 2020) and the final answer might not always follow the generated reasoning chain (Turpin et al., 2023). xLLM (Chuang et al., 2024) improves explanation fidelity from LLMs using a fidelity evaluator, tailored for classification tasks. We focus on explaining LLM generation by analyzing input prompt attributions to output content, independent of the model’s suboptimal reasoning abilities.
@@ 3 Preliminaries
There are limited prior works about explaining the relationship between the individual input prompts and the entire generated sequence, and their faithfulness is limited by treating input tokens independently. Specifically, Captum (Miglani et al., 2023), calculates the importance score for each token by slightly perturbing the input $\bm{x}$ at the $i$-th token:
@@ 4 Proposed Method
In order to explain language generation via prompt attribution, instead of treating the prompt tokens independently as in previous works (Miglani et al., 2023), we propose to evaluate the joint effect of $k$ prompt tokens on the generated sequence.
@@
Gradient-Guided Masking: Trivial solutions that set the mask $\bm{m}$ by uniformly random could cost massive sampling to hit the right optimization direction. Gradient is a common indicator of feature importance, as evidenced in existing practices (Ebrahimi et al., 2018; Zou et al., 2023; Shin et al., 2020). Therefore, we propose to use gradient as a guidance to set and sample the mask for more efficient optimization. Specifically, we begin with the binary mask $\bm{m}^{(0)}=\bm{1}$ which indicates that all tokens in the input $\bm{x}$ are marked as non-explanatory ones. Then we compute the gradient $\nabla_{\bm{m}^{(0)}}\mathcal{L}(\bm{m}^{(0)},\bm{x},\bm{y};\bm{\theta})$ of the loss function in Eq. (3), and denote the magnitude of gradients as $\bm{g}=|\nabla_{\bm{m}^{(0)}}\mathcal{L}(\bm{m}^{(0)},\bm{x},\bm{y};\bm{\theta})|$. Note that the components with larger gradient magnitudes in $\bm{g}$ imply that altering the corresponding tokens could result in a sharper change to the generated outputs. In order to initialize the explanatory $k$ tokens guided by the gradient magnitude, we set the binary mask $\bm{m}^{(1)}$ where $m^{(1)}_{i}=0$ indicates $g_{i}$ is the top-$k$ value in $\bm{g}$. The gradient guidance and mask initialization are obtained following Line 1-2 in the algorithm.
@@ 5.1 Experiment Settings
In the experiment, we employ two LLMs as our targeted $f_{\bm{\theta}}(\cdot)$: LlaMA-2 (7B-Chat) (Touvron et al., 2023) and Vicuna (7B) (Zheng et al., 2023). There are relatively few methods attributing prompts on the entire language generation, and we choose random removal (Random), Integrated-Gradient (Sundararajan et al., 2017), averaged attentions across all layers (Pruthi et al., 2019)(Attention), last-layer attention (Zhao and Shan, 2024)(Last-Attention) and Captum (Miglani et al., 2023) as our baseline for comparison. We also compare JoPA with ReAGent (Zhao and Shan, 2024), with the results presented in Appendix A.13. We implement these models using the PyTorch framework and pretrained weights from the transformers Python library (Wolf et al., 2020), and conduct our experiments on an Nvidia RTX A6000-48GB platform with CUDA version 12.0.
@@
We employ three distinct text generation datasets: Alpaca (Taori et al., 2023), tldr_news (Belvèze, 2022), and MHC (Amod, 2024), to evaluate the effectiveness of our method across various generation tasks. To capture the joint influence of prompts on model outputs, we use long-context prompts (15+ words) rather than simple sentences, as they convey more information and exhibit higher textual correlation. For evaluation, 110 samples are randomly selected from each publicly available dataset, detailed in AppendixA.2
@@ 5.2 Evaluation Metrics
Faithfulness scores are a key metric for assessing the quality of explanations, with faithful explanations accurately reflecting the model’s decision-making process (Jacovi and Goldberg, 2020). Studies (Samek et al., 2017; Hooker et al., 2019) suggest that if a certain input tokens are truly important, their removal should lead to a more significant change in model output than the removal of random tokens (Madsen et al., 2022). Therefore, after removing explanatory tokens identified by a faithful method, the model’s new generation would significantly differ from using the original output. Moreover, the input prompt with masking these explanatory tokens are less likely to reproduce the original model response $\bm{y}$, indicating a lower value of $p_{\bm{\theta}}(\bm{y}|\bm{m}\odot\bm{x})$.
@@
Word Frequency: BLEU (Papineni et al., 2002) evaluates how closely the candidate text $\bm{y}^{\prime}$ matches the reference $\bm{y}$ by measuring n-gram precision with clipping to avoid overcounting and brevity adjustments for shorter candidates. ROUGE-L (Lin, 2004) measures sequence overlap between $\bm{y}^{\prime}$ and $\bm{y}$, assessing precision, recall, and their F1 score.
Semantic Similarity: SentenceBert (Reimers and Gurevych, 2019) is a variation of the BERT model which is designed to generate high-quality sentence embeddings for the pairs of sentences. We leverage SentenceBert to transform the text $\bm{y}$ and $\bm{y}^{\prime}$ into fixed-length embedding vectors, and calculate the cosine similarity between their embeddings to quantify their semantic similarity.
Probability Measurement: We define two measurements to reflect how the likelihood of generating the original text $\bm{y}$ changes before and after applying the explanation mask. We first define the Probability Ratio (PR) to indicate how less likely to generate the original output $\bm{y}$ when explanatory tokens are masked: $\text{PR}(\bm{x},\bm{y},\bm{m})=\frac{\tilde{p}_{\bm{\theta}}(\bm{y}|\bm{m}\odot\bm{x})}{\tilde{p}_{\bm{\theta}}(\bm{y}|\bm{x})}$, where $\tilde{p}(\bm{y}|\bm{x})=p(\bm{y}|\bm{x})^{1/s}$ is the length-normalized generation probability. If the PR score is far below the random baseline, we can conclude that the masked tokens are indeed important to cause the model generating $\bm{y}$. In addition, we also calculate the KL-divergence between these two distributions to measure their difference: $D_{\text{KL}}(p_{\bm{\theta}}(\bm{y}|\bm{m}\odot\bm{x})||p_{\bm{\theta}}(\bm{y}|\bm{x}))$, and a larger score indicates a larger change in generation likelihood and a more accurate explanation.
@@ A.2 Dataset Details
  • Alpaca Taori et al. (2023) is a dataset with 52000 unique examples consisting of instructions and demonstrations generated by OpenAI’s text-davinci-003 engine. Each example in the dataset includes an instruction that describes the task the model should perform, accompanied by optional input context for that task. In our experiment, we randomly select a subset of these examples for verification purposes.
  • tldr_news Belvèze (2022) dataset is constructed by collecting a daily tech newsletter. For every piece of data, there is a headline and corresponding content extracted. The task is to ask the model to simplify the extracted content and then generate a headline from the input.
  • mental_health_counseling (MHC) Amod (2024) includes broad pairs of questions and answers derived from online counseling and therapy platforms. It covers a wide range of mental-health related questions and concerns, as well as the advice provided by the psychologists. It is utilized for used to fine-tuning the model to generate the metal health advice. Here, we prompt the model to generate an advice based on the provided question.
@@ A.3 Case Study
In this section, we visualize three examples which are selected from the dataset Alpaca, tldr_news, and MHC respectively, shown in Figure. 3. The tokens in the red frame are the $k$ explanatory tokens extracted by the model LlaMA-2 (7B-Chat), and the output responses framed in blue are the most informative results selected by human. In the first data sample, the recognized token: "brains" and "cognitive", has some semantic relationship intuitively. It is a common sense the cognitive ability is underpinned by the brain’s function and structure. There are numerous reports and studies focused on the functionally relationship between the brain and cognitive ability (Zhang, 2019). Due to the extensive amount of data on which the LLMs are trained, the model can integrate their relationship which is reflected in its response.